OpenStudy (anonymous):

A railroad tunnel is shaped like a semiellipse as shown below. The height of the tunnel at the center is 58 ft and the vertical clearance must be 29 ft at a point 21 ft from the center. Find an equation for the ellipse.

OpenStudy (anonymous):

Diagram: http://prntscr.com/7mjbvf

jimthompson5910 (jim_thompson5910):

not sure, but let me think

OpenStudy (anonymous):

Anything?

jimthompson5910 (jim_thompson5910):

is there anything about the ellipse not mentioned? I feel like there's some missing info

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

vertex is 0,58 and there are two points on the ellipse at +-21,29

jimthompson5910 (jim_thompson5910):

if the ellipse is taller than it is wide, then a = 58 and this pairs up with the y^2 term

jimthompson5910 (jim_thompson5910):

and if (21,29) is on the ellipse, then (x,y) = (21,29) --> x = 21 and y = 29

jimthompson5910 (jim_thompson5910):

assume the center is (h,k) = (0,0)

jimthompson5910 (jim_thompson5910):

\[\Large \frac{(x-h)^2}{b^2}+\frac{(y-k)^2}{a^2} = 1\] \[\Large \frac{(21-0)^2}{b^2}+\frac{(29-0)^2}{58^2} = 1\] solve for b

OpenStudy (anonymous):

24.25ish

jimthompson5910 (jim_thompson5910):

actually you don't have to solve for b you can stop at b^2

OpenStudy (anonymous):

well f***

jimthompson5910 (jim_thompson5910):

since b^2 is in the denominator under the x^2 term

OpenStudy (anonymous):

I'm getting a "Max Iterations Error" on my calculator

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (anonymous):

TI-36x Pro

jimthompson5910 (jim_thompson5910):

I'm not familiar with that type

jimthompson5910 (jim_thompson5910):

but you should get b^2 = 588

OpenStudy (anonymous):

It has a number solve function where I can type in 2 sides of an equation and It'll solve for a variable in the equation. That's what I was using and it gave that error. When I solved for b^2 it works

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

hmm strange

OpenStudy (anonymous):

ok so we found b^2 and b, so now what?

jimthompson5910 (jim_thompson5910):

replace b^2 with 588 and a^2 with whatever 58^2 is

jimthompson5910 (jim_thompson5910):

(h,k) = (0,0)

jimthompson5910 (jim_thompson5910):

x and y are left alone in the equation

OpenStudy (anonymous):

so \[\large \frac{ x^2 }{ 588 }+\frac{ y^2 }{ 3364 } =1\]

OpenStudy (anonymous):

then we need to add the values

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

add values?

OpenStudy (anonymous):

is that the equation or is there something left?

jimthompson5910 (jim_thompson5910):

the last thing you posted is the equation they want

jimthompson5910 (jim_thompson5910):

I guess you could solve for y to get some expression in the form y = ... that will get you the top half of the ellipse, which is the tunnel

OpenStudy (anonymous):

should I do that?

jimthompson5910 (jim_thompson5910):

hmm now I'm not sure if they want the full ellipse or just the upper half ellipse

OpenStudy (anonymous):

ill give both

jimthompson5910 (jim_thompson5910):

good idea

OpenStudy (anonymous):

So do I just solve for y=?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

So I'm stuck at \[\frac{ y^2 }{ 3364 }=1-\frac{ x^2 }{ 588 }\]

jimthompson5910 (jim_thompson5910):

multiply both sides by 3364 then take the square root of both sides you focus on the positive square root because you want the upper half

OpenStudy (anonymous):

how I multiply x^2/588 by 3364 and what would that come out to

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ y^2 }{ 3364 }=1-\frac{ x^2 }{ 588 }\] \[\Large 3364*\frac{ y^2 }{ 3364 }=3364*(1-\frac{ x^2 }{ 588 })\] \[\Large y^2=3364*(1-\frac{ x^2 }{ 588 })\] \[\Large y^2=3364*1-3364*\frac{ x^2 }{ 588 }\] \[\Large y^2=3364-\frac{3364x^2 }{ 588 }\] \[\Large y^2=3364-\frac{841x^2 }{ 147 }\] \[\Large y=???\]

OpenStudy (anonymous):

It was that last part I was confused about.\[y=58-\frac{ 29x }{ 7\sqrt{3} }\]

jimthompson5910 (jim_thompson5910):

you can't take the square root like that

OpenStudy (anonymous):

well crap

jimthompson5910 (jim_thompson5910):

\[\Large y^2=3364-\frac{841x^2 }{ 147 }\] \[\Large \sqrt{y^2}=\sqrt{3364-\frac{841x^2 }{ 147 }}\] \[\Large y=\sqrt{3364-\frac{841x^2 }{ 147 }}\]

jimthompson5910 (jim_thompson5910):

you apply the square root to the entire side

OpenStudy (anonymous):

then what? Is that all we can do

jimthompson5910 (jim_thompson5910):

yeah that's as far as you can go

OpenStudy (anonymous):

Great thanks

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{x + y} \ne \sqrt{x} + \sqrt{y}\]

OpenStudy (anonymous):

It's been a long day of math for me. Finally off to bed. Thanks for all the help, dude!

jimthompson5910 (jim_thompson5910):

no problem